A NEW COLLISION-INDUCED TRANSITION IN THE VACUUM ULTRAVIOLET SPECTRUM OF O_{2}

L. J. COIN, B. R. LEWIS, and \underline{S. T. GIBSON}, Research School of Physical Sciences and Engineering, The Australian National University, Canberra, ACT 0200, Australia.

The collision-induced photoabsorption spectrum of O_{2}, pressurised by He, is presented in the 1180–1340 Å-region. In addition to the collision-broadened wings of the dipole-allowed mixed Rydberg-valence transitions, we have found a series of diffuse peaks whose intensities depend linearly on the foreign-gas pressure. The observed peaks have been assigned as the (1,0)–(5,0) bands of the ΔΛ = 2 collision-induced 3pπ_{u} ^{3}Δ_{u} ← X^{3}Σ_{g}^{−} Rydberg transition, partly on the basis of comparisons with the photoabsorption spectrum of metastable O_{2}(^{1}Δ_{g}).^{a} The relative regularity of the spacings between the observed peaks supports the conclusions of Buenker and Peyerimhoff,^{b} in an ab initio study, that there is little interaction between the 3pπ_{u} ^{3}Δ_{u} Rydberg state and the A′ ^{3}Δ_{u} valence state, in contrast to the situation for the corresponding states of ^{3}Σ_{u}^{−} and ^{3}Π_{u} symmetries. Although the complexity of this spectral region and the diffuseness of the bands precludes any detailed line-profile analyses, it is clear that the integrated intensity of the 3pπ_{u} ^{3}Δ_{u} ← X^{3}Σ_{g}^{−} Rydberg system exceeds that of the collision-induced A′ ^{3}Δ_{u} ← X^{3}Σ_{g}^{−} valence system^{c} by at least an order of magnitude. Possible sources for this intensity are discussed.

aP. C. Hill, Ph.D. Thesis, Australian National University, (1991).

bR. J. Buenker and S. D. Peyerimhoff, Chem. Phys. Lett. 34, 225 (1975).

cV. I. Dianov-Klokov, Opt. Spectrosc. 21, 233 (1996).